# Can Watermarks Survive Translation? On the Cross-lingual Consistency of Text Watermark for Large Language Models

### v1

@@ 6.1 LLM Watermarking
Text watermarking aims to embed a watermark into a text and detect the watermark for any given text. Currently, text watermark method can be classified into two categories Liu et al. (2023c): watermarking for exsiting text and watermarking for generated text. In this work, we focus on the latter, which is more challenging and has more practical applications.
This type of watermark method usually can be illustrated as the watermark ironing process (modifying the logits of the LLM during text generation) and watermark detection process (assess the presence of watermark by a calculated watermark strength score).  Kirchenbauer et al. (2023a) introduces KGW, the first watermarking method for LLMs.  Hu et al. (2023) proposes UW without affecting the output probability distribution compared to KGW.  Liu et al. (2023b) introduces SIR, a watermarking method taking into account the semantic information of the text, which shows robustness to paraphrase attacks.  Liu et al. (2023a) proposes the first unforgeable publicly verifiable watermarking algorithm for large language models. SemStamp  Hou et al. (2023) is another semantic-related watermarking method and it generate watermarked text at sentence granularity instead of token granularity.  Tu et al. (2023) introduces WaterBench, the first comprehensive benchmark for LLM watermarks. We introduce the details of KGW, UW and SIR in § 2.2.
@@ 6.2 Watermark Robustness
A good watermarking method should be robust to various watermarking removal attacks. However, current works on watermarking robustness mainly focus on single-language attacks, such as paraphrase attacks. For example,  Kirchenbauer et al. (2023b) evaluates the robustness of KGW against paraphrase attacks as well as copy-paste attacks and proposes a detect trick to improve the robustness to copy-paste attacks.  Zhao et al. (2023) employs a fixed green list to improve the robustness of KGW against paraphrase attacks and editing attacks.  Chen et al. (2023) proposes a new paraphrase robust watermarking method “XMark” based on “text redundancy” of text watermark.
@@ 7 Conclusion
This work aims to investigate the cross-lingual consistency of watermarking methods for LLMs. We first characterize and evaluate the cross-lingual consistency of current watermarking techniques for LLMs, revealing that current watermarking methods struggle to maintain their watermark strengths across different languages. Based on this observation, we propose the cross-lingual watermark removal attack (CWRA), which significantly challenges watermark robustness by efficiently eliminating watermarks without compromising performance. Through the analysis of two primary factors that influence cross-lingual consistency, we propose X-SIR as a defense strategy against CWRA. Despite its limitations, this approach greatly improves the AUC and paves the way for future research. Overall, this work completes a closed loop in the study of cross-lingual consistency in watermarking, including: evaluation, attacking, analysis, and defensing.
@@ 8 Limitations
While X-SIR has demonstrated promising capabilities in enhancing cross-lingual consistency and defending against CWRA, it also faces certain limitations. A key limitation is its narrow scope of applicability. This limitation stems from the reliance on an external bilingual dictionary for semantic clustering of the vocabulary, which means X-SIR focuses solely on whole words and neglects smaller word units. Consequently, its effectiveness is closely tied to the tokenization. X-SIR’s performance may be compromised if the tokenizer favors finer-grained token segmentation. Moreover, X-SIR could face difficulties in scenarios with a significant difference in word order between the original and pivot languages. This aspect is crucial, as attackers can exploit these differences in any language pair to conduct CWRA. Therefore, X-SIR does not solve the issue of cross-lingual consistency but sets the stage for future research.
Appendix A Watermark Method Detail
In this section, we provide more details about the watermarking methods we use in our experiments.
A.1 Watermark Strength Score
@@
As we discussed in Section 6.1, we focus on the watermarking methods for LLMs. And we illustrated that a watermarking method can be divided into two parts: watermark ironing process and watermark detection process. In ironing process, the watermark is embedded into the text by modifying the logits of the LLM during text generation. In detection process, the watermark detector calculates the watermark strength score $S$ to assess the presence of watermark. $S$ is a scalar value to indicate the strength of the watermark in the text. For any given text, we can calculate its watermark strength score $S$ based on detection process of the watermarking method. A higher $S$ indicates that the text is more likely to contain watermark. In the opposite, a lower $S$ indicates that the text is less likely to contain watermark. Every watermarking method has its own way to ironing the watermark and calculate the watermark strength score $S$. We provide the details of watermark ironing process and watermark detection process for KGW, UW and SIR in the following sections.
A.2 KGW
@@
In Section 2.2, we introduce the watermark ironing process and watermark detection process of KGW. Here we provide the details of experiment settings for KGW. KGW uses a hash function $H$ to compute a hash of the previous $k$ tokens. In this work, we follow the experiment settings of  Kirchenbauer et al. (2023a), using the minhash with $k=4$. The proportion of green token lists $\mathcal{V}_{g}$ to the total word list $\mathcal{V}$ is set to $\gamma=0.25$. The constant bias $\delta$ is set to $2.0$.
A.3 UW
In this section, we introduce the detail watermarking ironing process and watermark detection process of UW. The ironing process of UW is similar to KGW. The difference between UW and KGW is the way to modify the logits:
@@
Then we get the adjusted logits $\tilde{P}_{M}(x^{n+1}|\boldsymbol{x}^{1:n})$.
@@
Where $r_{i}(x_{k})$ is the LLR of token $x_{k}$ at position $i$. To avoid the limitation of the original LLR,  Hu et al. (2023) proposes a maximin variant LLR. The calculating process of maximin variant LLR can be illustrated as follows:
@@
A.4 SIR
In this section, we introduce the watermark ironing process and watermark detection process of SIR. Given a language model $M$, a text embedding language model $E$, a trained watermark model $T$, previous generated text $\boldsymbol{t}=\left[t^{0},\cdots,t^{l-1}\right]$. The watermark ironing process of SIR can be illustrated as follows:
  1. (1)
    Generate the next token logits: $P_{M}(t^{l}|\boldsymbol{t})$.
  2. (2)
    Generate the semantic embedding of the previous tokens: $e_{l}=E(\boldsymbol{t})$.
  3. (3)
    Generate the watermark logit bias: $\Delta=T(e_{l})\in\mathbb{R}^{|\mathcal{V}|}$.
  4. (4)
    Adjust the logits: $\tilde{P}_{M}(t^{l}|\boldsymbol{t})=P_{M}(t^{l}|\boldsymbol{t})+\delta\times\Delta$.
Where $\delta$ is a hyperparameter (a small positive number) to control the strength of the watermark.
The detection process calculates the mean of the watermark bias of the detected text to assess the watermark strength score. Given the detected text $\boldsymbol{x}=\left[x^{1},\cdots,x^{N}\right]$.
@@ A.4 SIR
where $I(x^{n})$ is the index of token $x^{n}$ within the vocabulary $\mathcal{V}$. $\Delta_{I(x^{n})}(\boldsymbol{x}^{1:n-1})$ is the watermark bias of token $x^{n}$ at position $n$. Since the watermark bias satisfies the unbiased property: $\sum_{t\in\mathcal{V}}\Delta_{I(t)}=0$, the expected detection score for the text without watermark is 00. The detection score of the text with watermark should over 00.
In this work, we follow the origin experiment settings for the watermark model $T$. The watermark model uses a four-layer fully connected residual network with rectified linear unit activations, with the hyperparameter $k_{1}=20,k_{2}=1000,\lambda_{1}=10,\lambda_{2}=0.1$. And we use Adam optimizer with learning rate $51e-$ 5 to train the watermark model $T$. The watermark strength parameter $\delta$ is set to $4.0$.
Figure 7: Trends of watermark strengths with text length before and after translation. Top three subfigures show the trends of watermark strengths with text length before and after translation for Baichuan-7B. Bottom three subfigures show the trends of watermark strengths with text length before and after translation for Llama-2-7b-chat.
Appendix B Result of Cross-lingual Consistency
In this section, we provide the detail result of our cross-lingual consistency experiment. We perform our experiment on three watermarking methods: KGW, UW and SIR and two large language models: Baichuan-7B and Llama-2-7b-chat. Figure 7 shows the result of cross-lingual consistency experiment.
All watermarking methods show that the watermark strength score of raw text is positively correlated with the text length. After translation, the watermark strength score significantly decreases. The result indicates that the watermarking methods we use in our experiments exhibit a lack of cross-lingual consistency.
Appendix C Result of Watermark Removal Attack
In this section, we provide the detail result of text watermark removal attack experiment. We perform the experiment on three watermarking methods: KGW, UW and SIR and two tasks: text summarization and question answering. The large language model we use in the experiment is Baichuan-7B. Figure 8 shows the ROC curves result of text summarization task under various attack methods: Re-translation, Paraphrase and CWRA for KGW, UW and SIR. While Figure 9 shows the ROC curves result of question answering task under various attack methods: Re-translation, Paraphrase and CWRA for KGW, UW and SIR. We also report the TPR values at a fixed FPR of $0.1$ in each subfigure.
All the results show that CWRA can effectively remove the watermark from the watermarked text which is more effective than Re-translation and Paraphrase attack.
Figure 8: ROC curves for KGW, UW and SIR under various attack methods: Re-translation, Paraphrase and CWRA. We also present TPR values at a fixed FPR of $0.1$. This is the result of text summarization task.
Figure 9: ROC curves for KGW, UW and SIR under various attack methods: Re-translation, Paraphrase and CWRA. We also present TPR values at a fixed FPR of $0.1$. This is the result of question answering task.

### v2

@@ 6.1 LLM Watermarking
Text watermarking aims to embed a watermark into a text and detect the watermark for any given text. Currently, text watermark method can be classified into two categories Liu et al. (2023): watermarking for existing text and watermarking for generated text. In this work, we focus on the latter, which is more challenging and has more practical applications.
This type of watermark method usually can be illustrated as the watermark ironing process (modifying the logits of the LLM during text generation) and watermark detection process (assess the presence of watermark by a calculated watermark strength score). Kirchenbauer et al. (2023a) introduces KGW, the first watermarking method for LLMs. Hu et al. (2023) proposes UW without affecting the output probability distribution compared to KGW. Liu et al. (2024b) introduces SIR, a watermarking method taking into account the semantic information of the text, which shows robustness to text re-writing attacks. Liu et al. (2024a) proposes the first unforgeable and publicly verifiable watermarking algorithm for LLMs. SemStamp Hou et al. (2023) is another semantic-related watermarking method and it generate watermarked text at sentence granularity instead of token granularity. Tu et al. (2023) introduces WaterBench, the first comprehensive benchmark for LLM watermarks.
@@ 6.2 Watermark Robustness
A good watermarking method should be robust to various watermarking removal attacks. However, current works on watermarking robustness mainly focus on single-language attacks, such as paraphrase attacks. For example, Kirchenbauer et al. (2023b) evaluates the robustness of KGW against paraphrase attacks as well as copy-paste attacks and proposes a detect trick to improve the robustness to copy-paste attacks. Zhao et al. (2023) employs a fixed green list to improve the robustness of KGW against paraphrase attacks and editing attacks. Chen et al. (2023) proposes a new paraphrase robust watermarking method “XMark” based on “text redundancy” of text watermark. Lu et al. (2024) proposes an entropy-based text watermarking detection method that achieves better detection performance in low-entropy scenarios.
@@ 7 Conclusion
This work aims to investigate the cross-lingual consistency of watermarking methods for LLMs. We first characterize and evaluate the cross-lingual consistency of current watermarking techniques, revealing that current watermarking methods struggle to maintain their watermark strengths across different languages. Based on this observation, we propose the cross-lingual watermark removal attack (CWRA), which significantly challenges watermark robustness by efficiently eliminating watermarks without compromising text quality. Through the analysis of two primary factors that influence cross-lingual consistency, we propose X-SIR as a defense strategy against CWRA. Despite its limitations, this approach greatly improves watermark detection performance under cross-lingual scenario and paves the way for future research. Overall, this work completes a closed loop in the study of cross-lingual consistency in watermarking, including: evaluation, attacking, analysis, and defensing.
Since our first release, the X-SIR algorithm has been integrated into MarkLLM Pan et al. (2024), an open-source toolkit for LLM Watermarking.
@@ 8 Limitations
X-SIR relies on the semantic clustering described in Algorithm 1 which only considers tokens shared by the vocabulary $\mathcal{V}$ of the model and the external dictionary $\mathcal{D}$. This design results in the following limitations:
  • Language coverage: X-SIR only supports languages supported by the model. However, in a real scenario, an attacker can choose the original and the pivot language at will. This limitation has revealed by Figure 6, where X-SIR brings a much more significant boost to Ja & Zh than to De & Fr since the vocabulary of the LLM (Baichuan-7B) supports Zh & Ja better.
  • Vocab coverage: Since the external dictionary $\mathcal{D}$ only contains whole words, word units can not be clustered in Algorithm 1 and will left as isolated nodes. Consequently, X-SIR’s performance might be compromised if the tokenizer favors finer-grained token segmentation.
X-SIR does not solve the issue of cross-lingual consistency but sets the stage for future research.
Acknowledgements
This paper is mainly supported by the Tencent AI Lab Fund (RBFR2024002). Zhiwei and Rui are partially supported by the National Natural Science Foundation of China (62176153) and the Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102, as the MoE Key Lab of Artificial Intelligence, AI Institute, Shanghai Jiao Tong University). Zhuosheng is partially supported by the Joint Funds of the National Natural Science Foundation of China (Grant No. U21B2020).
Appendix A Details of Watermarking Methods
@@
As discussed in § 6.1, we focus on the watermarking methods for large language models (LLMs). A watermarking method can be divided into two processes: the watermark ironing process and the watermark detection process. In ironing process, the watermark is embedded into the text by modifying the logits of the LLM during text generation. In detection process, the watermark detector calculates the watermark strength score $s$ to assess the presence of watermark. $s$ is a scalar value to indicate the strength of the watermark in the text. For any given text, we can calculate its watermark strength score $s$ based on detection process of the watermarking method. A higher $s$ indicates that the text is more likely to contain watermark. In the opposite, a lower $s$ indicates that the text is less likely to contain watermark. Every watermarking method has its own way to ironing the watermark and calculate the watermark strength score $s$. We detail KGW, UW and SIR in the following sections.
A.1 KGW
@@
In § 2.2, we introduce the processes of watermark ironing and watermark detection in the KGW method. Here, we detail the experimental settings employed for KGW. KGW uses a hash function $H$ to compute the hash of the previous $k$ tokens. In this work, we adhere to the experimental setting reported by Kirchenbauer et al. (2023a), employing the hash function minhash with $k=4$. The ratio of green token lists $\mathcal{V}_{g}$ to the total word list $\mathcal{V}$ is set at $\gamma=0.25$. Additionally, the constant bias $\delta$ is fixed at $2.0$.
A.2 UW
The ironing process in UW is analogous to that in KGW; however, the two differ in their respective function of modifying the logits. Here is the detail watermark ironing process in UW:
@@
Then we get the adjusted probability of next token $\tilde{P}_{M}(x^{n+1}|\boldsymbol{x}^{1:n})$.
@@
Where $r_{i}(x_{k})$ is the LLR of token $x_{k}$ at position $i$. UW use a maximin variant LLR  Hu et al. (2023) to avoid the limitation of the origin LLR. The calculating process of maximin variant LLR can be formulated as follows:
@@
A.3 SIR
As introduced in § 2.2, the ironing process in SIR assigns a watermark bias, $\Delta_{i}$, to every token $v_{i}$.
For any given text, the watermark detector calculates the mean watermark bias to determine the watermark strength score. Consider the detected text represented as $\boldsymbol{x}=\left(x^{1},\ldots,x^{N}\right)$. The watermark strength score, $s$, can be expressed by the following equation:
@@ A.3 SIR
where $I(x^{n})$ indicates the index of the token $x^{n}$ within the vocabulary $\mathcal{V}$, and $\Delta_{I(x^{n})}(\boldsymbol{x}^{1:n-1})$ denotes the watermark bias for token $x^{n}$ at position $n$. Given that the watermark bias satisfies the unbiased property, $\sum_{t\in\mathcal{V}}\Delta_{I(t)}=0$, the expected detection score for normal text is 00. Consequently, the detection score for watermarked text should exceed 00 significantly.
Appendix B Full Result of Watermark Removal Attack
Figure 9 presents the full results of watermark removal attacks. It is equivalent to Figure 4, but presents results on text summarization and question answering separately.
(a) Text summarization
(b) Question answering
Figure 9: ROC curves for KGW, UW and SIR under various attack methods: Re-translation, Paraphrase and CWRA. We also present TPR values at a fixed FPR of $0.1$.
Appendix C Full Result of X-SIR
Figure 10 presents full watermark detection performance (ROC curves, AUCs and TPRs) of X-SIR. We tested it on the following LLMs:
We tested it under the following attack methods:
  • No attack
  • Paraphrase attack (§ 4.1)
[…]
We also plot the performance of SIR as baselines.
(a) Baichuan-7B
(b) Baichuan2-7B-Base
(c) Llama-2-7B
(d) Mistral-7B-v0.1
Figure 10: ROC curves for X-SIR and SIR under various attack methods: no attack, paraphrase and direct translation. We also present AUC and TPR values at a fixed FPR of 0.1.